# Avoiding Catastrophe in Online Learning by Asking for Help

### v5

@@ B.3 Adaptive adversaries
If $s_{t}$ is allowed to depend on the events of prior time steps, we say that the adversary is adaptive. In contrast, a non-adaptive or “oblivious” adversary must choose the entire input upfront. This distinction is not relevant for deterministic algorithms, since an adversary knows exactly how the algorithm will behave for any input. In other words, the adversary gains no new information during the execution of the algorithm. For randomized algorithms, an adaptive adversary can base the choice of $s_{t}$ on the results of randomization on previous time steps (but not on the current time step), while an oblivious adversary cannot.
In the standard online learning model, Hedge guarantees sublinear regret against both oblivious and adaptive adversaries (Chapter 5 of Slivkins et al. (2019) or Chapter 21 of Shalev-Shwartz & Ben-David (2014)). However, Russo et al. (2024) state their result only for oblivious adversaries. In order for our overall proof of Theorem 5.1 to hold for adaptive adversaries, Lemma 5.4 (Lemma 3.5 in Russo et al., 2024) must also hold for adaptive adversaries. In this section, we argue why the proof of Lemma 5.4 (Lemma 3.5 in their paper) goes through for adaptive adversaries as well. For this rest of Section B.3, lemma numbers refer to the numbering in Russo et al. (2024).
@@
Recall from Section B.1 that Russo et al. (2024) allow two separate parameters $k$ and $\hat{k}$, which we unify for simplicity. Recall also that Lemma 3.5 refers to the variant of Hedge which queries with probability $p=\hat{k}/T=k/T$ independently on each time step (Algorithm 2. More precisely, on each time step $t$, the algorithm samples $X_{t}\sim\text{Bernoulli}(p)$ and queries if $X_{t}=1$. The key idea is that $X_{t}$ is independent of events on previous time steps. Thus even conditioning on the history up to time $t$, for any for any random variable $Y_{t}$ we can write
@@
The other tricky part of the proof is applying Observation 3.3 using a new loss function $\hat{\ell}$ defined by ^$\hat{\ell}_{t}(i)=\frac{T}{\hat{k}}(\ell_{t}(i)-\ell_{t}(i_{t}^{*}))\mathbf{1}(X_{t}=1)$. To do so, we must argue that standard Hedge run on $\hat{\ell}$ is the “counterpart without queries” of HedgeWithQueries. Specifically, both algorithms must have the same weight vectors on every time step, and the only difference should be that HedgeWithQueries takes the optimal action on each time step independently with probability $p$ (and otherwise behaves the same as standard Hedge). On time steps with $X_{t}=0$, standard Hedge observes ^$\hat{\ell}_{t}(i)=0$ for all actions $i$ and thus makes no updates, and HedgeWithQueries makes no updates by definition. On time steps with $X_{t}=1$, both algorithms perform the typical updates $w_{t+1}(i)=w_{t}(i)\cdot\exp(-\eta(\hat{\ell}_{t}(i)-\hat{\ell}_{t}(i_{t}^{*})))$. Thus the weight vectors are the same for both algorithms on every time step. Furthermore, HedgeWithQueries takes the optimal action at time $t$ iff $X_{t}=1$, which occurs independently with probability $p$ on each time step. Thus standard Hedge run on $\hat{\ell}$ is the “counterpart without queries” of HedgeWithQueries. Note that since $\hat{\ell}$ is itself a random variable, the law of iterated expectation is necessary to formalize this.
@@
We use the standard “one versus rest” reduction (see, e.g., Chapter 29 of Shalev-Shwartz & Ben-David, 2014). For each action $y$, we will learn a binary classifier which predicts whether action $y$ is the mentor’s action. Formally, for each $y\in\mathcal{Y}$, define the policy class $\Pi_{y}=\{\pi_{y}:\pi\in\Pi\text{ and }\pi_{y}(x)=\mathbf{1}(\pi(x)=y))\ \forall x\in\mathcal{X}\}$. In words, for each policy $\pi:\mathcal{X}\to\mathcal{Y}$ in $\Pi$, there exists a policy $\pi_{y}:\mathcal{X}\to\{0,1\}$ in $\Pi_{y}$ such that $\pi_{y}(x)=\mathbf{1}(\pi(x)=y)$ for all $x\in\mathcal{X}$.
@@
        $\tilde{\Pi}_{y}\leftarrow$ any smooth $\varepsilon$-cover of $\Pi$ of size at most $(41/\varepsilon)^{d}$
     else if $\Pi$ has Littlestone dimension $d$ then
        $\tilde{\Pi}_{y}\leftarrow$ any adversarial cover of size at most $(eT/d)^{d}$
  for $t$ from $1$ to $T$ do
@@
    Let $\pi^{my}:\mathcal{X}\to\{0,1\}$ be the policy defined by $\pi^{my}(x)=\mathbf{1}(\pi^{m}(x_{t})=y)$. Note that querying the mentor tells the agent $\pi^{m}(x_{t})$, which allows the agent to compute $\pi^{my}(x_{t})$: this is necessary when Algorithm 1 queries while running on some $\Pi_{y}$.
@@
The argument is similar to the proof of Lemma B.2. If $\mu_{t}^{m}(x_{t})\neq\mu_{t}(x_{t},y_{t})$, then $y_{t}=y$ for some $y\in\mathcal{Y}$ where $b_{t}^{y}=1$. Therefore copy $y$ of Algorithm 1 did not query at time $t$ and $\pi_{t}^{y}(x_{t})=1$. Let $(x^{\prime},y^{\prime})=\operatorname*{arg\,min}_{(x,y)\in S_{t}^{y}:\pi_{t}^{y}(x_{t})=y}||x_{t}-x||$. Then $||x_{t}-x^{\prime}||\leq\varepsilon^{1/n}$ and $y^{\prime}=\pi_{t}^{y}(x_{t})=1$.
@@ Appendix D There exist policy classes which are learnable in our setting but not in the standard online model
Choosing $g(T)=T^{c}$ for $c\in(1/2,1)$ is sufficient to subconstant regret and sublinear queries:
@@ D.2 Proof notation
    Let $\mathcal{B}_{V}^{\prime}=\{B\in\mathcal{B}_{V}:\text{no descendant of $B$ is in $\mathcal{B}_{V}$}\}$.
@@ D.4 Proof
Since depth 0 buckets overlap only at their boundaries, and each depth 0 bucket has length $1/g(T)$, at most $g(T)\max_{t,t^{\prime}\in[T]}|x_{t}-x_{t^{\prime}}|=g(T)\operatorname*{diam}(\boldsymbol{x})$ depth 0 buckets are subsets of the interval $[\min_{t\in[T]}x_{t},\max_{t\in[T]}x_{t}]$. At most two depth 0 buckets are not subsets of that interval (one at each end), so the total number of depth 0 buckets is at most $g(T)\operatorname*{diam}(\boldsymbol{x})+2$.
@@
If we query at time $t$, then $\mu_{t}(x_{t},y_{t})=\mu_{t}^{m}(x_{t})$. Thus assume we do not query at time $t$: then there exists $x^{\prime}\in B_{t}$ (as defined on line 23 of Algorithm 4) such that $y_{t}=\pi(x^{\prime})=\pi^{m}(x^{\prime})$. Since $x_{t}$ and $x^{\prime}$ are both in $B_{t}$, $|x_{t}-x^{\prime}|\leq\operatorname*{len}(B_{t})$. Then by local generalization, $\mu_{t}(x_{t},y_{t})=\mu_{t}(x_{t},\pi^{m}(x^{\prime}))\geq\mu_{t}^{m}(x_{t})-L||x_{t}-x^{\prime}||\geq\mu_{t}^{m}(x_{t})-L\operatorname*{len}(B_{t})$. ∎
@@
Now consider some $B^{\prime}\in\mathcal{B}_{V}^{\prime}$ with $B\neq B^{\prime}$. We $|B\cap B^{\prime}|\leq 1$: otherwise one must be the descendant of the other, which contradicts the definition of $\mathcal{B}_{V}^{\prime}$. Suppose $B^{\prime}$ also intersects both $X_{j}$ and $X_{j+1}$: since $B^{\prime}$ is also an interval, $B^{\prime}$ must also contain an open neighborhood centered on the boundary between those two segments. But then $|B\cap B^{\prime}|>1$, which is a contradiction.
Therefore any pair of adjacent segments $X_{j}$ and $X_{j+1}$, there is at most one bucket in $\mathcal{B}_{V}^{\prime}$ which contains an open neighborhood around their boundary. Since there are at most $K-1$ pairs of adjacent segments, we have $|\mathcal{B}_{V}^{\prime}|\leq K-1\leq K$.∎
@@ Appendix E Properties of local generalization
Assume that $\mu$ satisfies Lipschitz continuity: for all $x,x^{\prime}\in\mathcal{X}$ and $y\in\mathcal{Y}$, $|\mu(x,y)-\mu(x^{\prime},y)|\leq L||x-x^{\prime}||$. Also assume that $\mu(x,\pi^{m}(x))=\max_{y\in\mathcal{Y}}\mu(x,y)$ for all $x\in\mathcal{X}$. Then $\mu$ satisfies local generalization with constant $2L$.
@@
$\mu(x,\pi^{m}(x^{\prime}))\geq$ $\ \mu(x^{\prime},\pi^{m}(x^{\prime}))-L||x-x^{\prime}||$ (Lipschitz continuity of $\mu$)
$\geq$ $\ \mu(x^{\prime},\pi^{m}(x))-L||x-x^{\prime}||$ ($\pi^{m}$ is optimal for $x^{\prime}$)
$\geq$ $\ \mu(x,\pi^{m}(x))-2L||x-x^{\prime}||$ (Lipschitz continuity of $\mu$ again)
$=$ $\ \mu^{m}(x)-2L||x-x^{\prime}||$ (Definition of $\mu^{m}(x)$)
@@
$\mu^{m}(x)+2L||x-x^{\prime}||\geq\mu^{m}(x)\geq\mu(x,\pi^{m}(x^{\prime}))$
So $-2L||x-x^{\prime}||\leq\mu(x,\pi^{m}(x^{\prime}))-\mu^{m}(x)\leq 2L||x-x^{\prime}||$ and therefore $|\mu(x,\pi^{m}(x^{\prime}))-\mu^{m}(x)|\leq 2L||x-x^{\prime}||$.∎
@@
Define $f(T):=\lceil\sqrt{g(T)T}\rceil$; by Lemma A.2, $g(T)\in o(f(T))$ and $f(T)\in o(T)$. Divide $\mathcal{X}$ into $f(T)$ equally sized sections $X_{1},\dots,X_{f(T)}$ in the exactly the same way as in Section 4.2; see also Figure 1. Assume that each $x_{t}$ is in exactly one section: this assumption holds with probability 1, so it does not affect the regret.
@@
Now fix $i$ and assume $j_{r}\neq j^{m}\ \forall r<i$. Queries in sections other than $j^{m}$ provide no information about the value of $j^{m}$, so $j^{m}$ is uniformly distributed across the set of sections not yet queried, i.e., $\{j\in[f(T)]:j_{r}\neq j\ \forall r<i\}$. There are at least $f(T)-i+1$ such sections, since there are $i-1$ prior queries at this point. Thus $\Pr[j_{i}\neq j^{m}\mid j_{r}\neq j^{m}\ \forall r<i]\geq\frac{f(T)-i}{f(T)-i+1}$ (the inequality is because this probability could also be 1 if $j_{i}=j_{r}$ for some $i<r$). Therefore
@@
Since $\varepsilon\in(0,1]$, we have $01-\log\varepsilon>$ 0. Also, $\Pr[\prod_{t=1}^{T}\mu(x_{t},y_{t})\leq\varepsilon]\geq\Pr[\prod_{t=1}^{T}\mu(x_{t},y_{t})=0]$, so

### v6

@@ B.3 Adaptive adversaries
If $x_{t}$ is allowed to depend on the events of prior time steps, we say that the adversary is adaptive. In contrast, a non-adaptive or “oblivious” adversary must choose the entire input upfront. This distinction is not relevant for deterministic algorithms, since an adversary knows exactly how the algorithm will behave for any input. In other words, the adversary gains no new information during the execution of the algorithm. For randomized algorithms, an adaptive adversary can base the choice of $x_{t}$ on the results of randomization on previous time steps (but not on the current time step), while an oblivious adversary cannot.
In the standard online learning model, Hedge guarantees sublinear regret against both oblivious and adaptive adversaries (Chapter 5 of Slivkins et al. (2019) or Chapter 21 of Shalev-Shwartz & Ben-David (2014)). However, Russo et al. (2024) state their result only for oblivious adversaries. In order for our overall proof of Theorem 5.1 to hold for adaptive adversaries, Lemma 5.4 (Lemma 3.5 in Russo et al., 2024) must also hold for adaptive adversaries. In this section, we argue why the proof of Lemma 5.4 (Lemma 3.5 in their paper) goes through for adaptive adversaries as well. For the rest of Section B.3, lemma numbers refer to the numbering in Russo et al. (2024).
@@
Recall from Section B.1 that Russo et al. (2024) allow two separate parameters $k$ and $\hat{k}$, which we unify for simplicity. Recall also that Lemma 3.5 refers to the variant of Hedge which queries with probability $p=\hat{k}/T=k/T$ independently on each time step (Algorithm 2). More precisely, on each time step $t$, the algorithm samples $X_{t}\sim\text{Bernoulli}(p)$ and queries if $X_{t}=1$. The key idea is that $X_{t}$ is independent of events on previous time steps. Thus even conditioning on the history up to time $t$, for any random variable $Y_{t}$ we can write
@@
The other tricky part of the proof is applying Observation 3.3 using a new loss function $\hat{\ell}$ defined by $\hat{\ell}_{t}(i)=\frac{T}{\hat{k}}(\ell_{t}(i)-\ell_{t}(i_{t}^{*}))\mathbf{1}(X_{t}=1)$. To do so, we must argue that standard Hedge run on $\hat{\ell}$ is the “counterpart without queries” of HedgeWithQueries. Specifically, both algorithms must have the same weight vectors on every time step, and the only difference should be that HedgeWithQueries takes the optimal action on each time step independently with probability $p$ (and otherwise behaves the same as standard Hedge). On time steps with $X_{t}=0$, standard Hedge observes $\hat{\ell}_{t}(i)=0$ for all actions $i$ and thus makes no updates, and HedgeWithQueries makes no updates by definition. On time steps with $X_{t}=1$, both algorithms perform the typical updates $w_{t+1}(i)=w_{t}(i)\cdot\exp(-\eta(\hat{\ell}_{t}(i)-\hat{\ell}_{t}(i_{t}^{*})))$. Thus the weight vectors are the same for both algorithms on every time step. Furthermore, HedgeWithQueries takes the optimal action at time $t$ iff $X_{t}=1$, which occurs independently with probability $p$ on each time step. Thus standard Hedge run on $\hat{\ell}$ is the “counterpart without queries” of HedgeWithQueries. Note that since $\hat{\ell}$ is itself a random variable, the law of iterated expectation is necessary to formalize this.
@@
We use the standard “one versus rest” reduction (see, e.g., Chapter 29 of Shalev-Shwartz & Ben-David, 2014). For each action $y$, we will learn a binary classifier which predicts whether action $y$ is the mentor’s action. Formally, for each $y\in\mathcal{Y}$, define the policy class $\Pi_{y}=\{\pi_{y}:\pi\in\Pi\text{ and }\pi_{y}(x)=\mathbf{1}(\pi(x)=y)\ \forall x\in\mathcal{X}\}$. In words, for each policy $\pi:\mathcal{X}\to\mathcal{Y}$ in $\Pi$, there exists a policy $\pi_{y}:\mathcal{X}\to\{0,1\}$ in $\Pi_{y}$ such that $\pi_{y}(x)=\mathbf{1}(\pi(x)=y)$ for all $x\in\mathcal{X}$.
@@
   $\tilde{\Pi}_{y}\leftarrow$ any smooth $\varepsilon$-cover of $\Pi_{y}$ of size at most $(41/\varepsilon)^{d}$
  else if $\Pi_{y}$ has Littlestone dimension $d$ then
   $\tilde{\Pi}_{y}\leftarrow$ any adversarial cover of $\Pi_{y}$ of size at most $(eT/d)^{d}$
for $t$ from 11 to $T$ do
@@
    Let $\pi^{my}:\mathcal{X}\to\{0,1\}$ be the policy defined by $\pi^{my}(x)=\mathbf{1}(\pi^{m}(x)=y)$. Note that querying the mentor tells the agent $\pi^{m}(x_{t})$, which allows the agent to compute $\pi^{my}(x_{t})$: this is necessary when Algorithm 1 queries while running on some $\Pi_{y}$.
@@
[…]
@@ Appendix D There exist policy classes which are learnable in our setting but not in the standard online model
Choosing $g(T)=T^{c}$ for $c\in(1/2,1)$ is sufficient for subconstant regret and sublinear queries:
@@ D.2 Proof notation
[…]
@@ D.4 Proof
Since depth 0 buckets overlap only at their boundaries, and each depth 0 bucket has length $1/g(T)$, at most $g(T)\max_{t,t^{\prime}\in[T]}|x_{t}-x_{t^{\prime}}|=g(T)\operatorname*{diam}(\boldsymbol{x})$ depth 0 buckets are subsets of the interval $\min_{t\in[T]}x_{t},\max_{t\in[T]}x_{t}]$. At most two depth 0 buckets are not subsets of that interval (one at each end), so the total number of depth 0 buckets is at most $g(T)\operatorname*{diam}(\boldsymbol{x})+2$.
@@
[…]
@@
Now consider some $B^{\prime}\in\mathcal{B}_{V}^{\prime}$ with $B\neq B^{\prime}$. We have $|B\cap B^{\prime}|\leq 1$: otherwise one must be the descendant of the other, which contradicts the definition of $\mathcal{B}_{V}^{\prime}$. Suppose $B^{\prime}$ also intersects both $X_{j}$ and $X_{j+1}$: since $B^{\prime}$ is also an interval, $B^{\prime}$ must also contain an open neighborhood centered on the boundary between those two segments. But then $|B\cap B^{\prime}|>1$, which is a contradiction.
Therefore for any pair of adjacent segments $X_{j}$ and $X_{j+1}$, there is at most one bucket in $\mathcal{B}_{V}^{\prime}$ which contains an open neighborhood around their boundary. Since there are at most K1K-1 pairs of adjacent segments, we have $|\mathcal{B}_{V}^{\prime}|\leq K-1\leq K$.∎
@@ Appendix E Properties of local generalization
[…]
@@
[…]
@@
[…]
@@
Define $f(T):=\lceil\sqrt{g(T)T}\rceil$; by Lemma A.2, $g(T)\in o(f(T))$ and $f(T)\in o(T)$. Divide $\mathcal{X}$ into $f(T)$ equally sized sections $X_{1},\dots,X_{f(T)}$ in exactly the same way as in Section 4.2; see also Figure 1. Assume that each $x_{t}$ is in exactly one section: this assumption holds with probability 1, so it does not affect the regret.
@@
Now fix $i$ and assume $j_{r}\neq j^{m}\ \forall r<i$. Queries in sections other than $j^{m}$ provide no information about the value of $j^{m}$, so $j^{m}$ is uniformly distributed across the set of sections not yet queried, i.e., $\{j\in[f(T)]:j_{r}\neq j\ \forall r<i\}$. There are at least f(T)i+1f(T)-i+1 such sections, since there are i1i-1 prior queries at this point. Thus $\Pr[j_{i}\neq j^{m}\mid j_{r}\neq j^{m}\ \forall r<i]\geq\frac{f(T)-i}{f(T)-i+1}$ (the inequality is because this probability could also be 1 if $j_{i}=j_{r}$ for some $i<r$). Therefore
@@
Since $\varepsilon\in(0,1]$, we have $1-\log\varepsilon>0$. Also, $\Pr[\prod_{t=1}^{T}\mu(x_{t},y_{t})\leq\varepsilon]\geq\Pr[\prod_{t=1}^{T}\mu(x_{t},y_{t})=0]$, so